OpenStudy (anonymous):

If a series converges absolutely, does that impact its interval of convergence?

OpenStudy (anonymous):

I have a question that is asking me to prove that a series converges for all x, but it has negative terms so I can only prove absolute convergence. So that is why I am asking this question

OpenStudy (anonymous):

absolute convergence is stronger than conditional convergence

OpenStudy (anonymous):

so yes

OpenStudy (anonymous):

We didn't really get into all of the details of what conditional convergence was today

OpenStudy (anonymous):

Is it possible for a series to converge for all x if it converges absolutely?

OpenStudy (anonymous):

means not absolute convergence simple example \[\sum\frac{1}{n}\] does not converge but \[\sum(-1)^n\frac{1}{n}\] does

OpenStudy (anonymous):

not sure what you mean exactly you got an example?

OpenStudy (anonymous):

Yes. I am supposed to use a comparison test to show that \[\sum_{n=0}^{\infty}\frac{ x ^{3n} }{ 2n!+1 }\] converges for all x

OpenStudy (anonymous):

the terms look positive to me

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

x could be negative

OpenStudy (anonymous):

not sure how you use a comparison test for series, usually use the ratio test

OpenStudy (anonymous):

I think that counts

OpenStudy (anonymous):

But my point is that x could be negative

OpenStudy (anonymous):

in any case it is clear that the radius of convergence is \(\infty\)

OpenStudy (anonymous):

if x is negative, the terms could be negative

OpenStudy (anonymous):

oh of course x could be negative, but it is a fixed number, it does not change say \(x=-10\) for example

OpenStudy (anonymous):

If x is negative and n is odd, terms are negative

OpenStudy (anonymous):

the idea is to compute \[\lim_{n\to \infty}|\frac{a_{n+1}}{a_n}|\]

OpenStudy (anonymous):

then figure out which \(x\) values work, i.e make sure that ratio is less than one

OpenStudy (anonymous):

Yes, but you can't use the ratio test in the first place if there are negative terms... or so I thought. Then you can only use it for absolute convergence

OpenStudy (anonymous):

ok i see you are a bid confused here n goes to infinity, x is a variable, but it is fixed as in \(f(x)=\frac{1}{1-x}\) which is equal to \(f(x)=\sum x^n\) if \(-1<x<1\)

OpenStudy (anonymous):

in that example x could be negative, for example it could be \(-\frac{1}{2}\)

OpenStudy (anonymous):

But 1/(1-x) didn't have a stipulation against negative terms

OpenStudy (anonymous):

when you use the ratio test you are taking the absolute value of the terms, computing the limit, then deciding for which values of x the limit exists

OpenStudy (anonymous):

But my teacher said that you can't use the ratio test in the first place if you can have negative terms T_T

OpenStudy (anonymous):

ok lets take it slow you are confusing the ratio test for seeing if a sum converges, like say \[\sum\frac{1}{n^2}\]

OpenStudy (anonymous):

with checking for the radius of convergence of \[\sum_{n=1}^{\infty}\frac{x^n}{n^2}\]

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Although I can't find any part of my book saying anything against negative terms for the ratio test, but it does talk about absolute convergence

OpenStudy (anonymous):

i see to find the radius of convergence of \[\sum c_nx^n\] put \(a_n=c_nx^n\) and compute \[\lim_{n\to \infty}|\frac{a_{n+1}}{a_n}|\]

OpenStudy (anonymous):

your answer will have an \(x\) in it

OpenStudy (anonymous):

So there are NO stipulations on the interval of convergence if it converges absolutely?

OpenStudy (anonymous):

then make sure \[\lim_{n\to \infty}|\frac{a_{n+1}}{a_n}|<1\]which will depend on \(x\) that is the radius of convergence

OpenStudy (anonymous):

if what converges absolutely?

OpenStudy (anonymous):

If the series does

OpenStudy (anonymous):

It says that the ratio test is a tool to check for absolute convergence

OpenStudy (anonymous):

i think i an unable to explain this, it seems you have confused a sum, a number, like \[\sum_{n=1}^{\infty}\frac{1}{2^n}\] with a series that looks like \[f(x)=\sum_{n=1}^{\infty}x^n\]

OpenStudy (anonymous):

I just want to know that if it is possible for something to converge absolutely and converge for all x